OpenStudy (anonymous):

Which of the following is equivalent to log 37 6 rounded to three decimal places?

jimthompson5910 (jim_thompson5910):

Is the problem \[\Large \log_{37}(6)\] ???

OpenStudy (anonymous):

Yes but the 6 isn't (6) just 6

jimthompson5910 (jim_thompson5910):

so it's \[\Large \log_{37} 6\]

jimthompson5910 (jim_thompson5910):

that's basically the same thing

OpenStudy (anonymous):

hey katie you can use this site for direct calculation http://www.wolframalpha.com

OpenStudy (anonymous):

just write there log6/log37

jimthompson5910 (jim_thompson5910):

\[\Large \log_{37}(6) = \frac{\ln(6)}{\ln(37)}\] \[\Large \log_{37}(6) \approx 0.49620609290339\] \[\Large \log_{37}(6) \approx 0.496\] Use a calculator on step 2

OpenStudy (anonymous):

Idk how to calculate that with a calculator

jimthompson5910 (jim_thompson5910):

type ln(6)/ln(37) into the calculator

OpenStudy (anonymous):

just press log and than 6

OpenStudy (anonymous):

divide it by pressing log and then 37